Determine the point estimate of the population proportion, the margin of error for the following confidence interval and the individuals in the sample with the specified characteristic, X. for the sample size provided.

Lower bound = 0.146, upper bound= 0.634 n=1000

Point estimate of the population proportion=0.390

margin or error = 0.244

the number of individuals in the sample with the specified characteristic is?

(round to the nearest integer as needed)
Answer #1

Given that, sample size ( n ) = 1000

Lower bound = 0.146, Upper bound = 0.634

Point estimate of the population proportion ( p ) is

p = ( 0.634 + 0.146 )/2 = 0.390

margin of error ( E ) is,

E = ( 0.634 - 0.146)/2 = 0.244

The number of individuals in the sample with the specified characteristic is,

X = n * p = 1000 * 0.390 = 390

=> X = 390